# TALC: Time-Aligned Captions for Multi-Scene Text-to-Video Generation

### v2

@@ 3.1 TALC: Time-Aligned Captions for Multi-Scene T2V Generation
Here, $\tau$ represents the timestamp in the diffusion modeling setup, which is applied during training as well as inference. We illustrate the framework in Figure 3. While TALC aims to equip the generative model with the ability to depict all the events in the multi-scene descriptions, the visual consistency is ensured by the temporal modules (attentions and convolution blocks) in the denoiser network. By design, our approach can be applied to the pretrained T2V model during inference.
3.2 Baselines
Here, we describe the baseline methods that could be used to generate videos for the multi-scene descriptions from a given diffusion text-to-video generative model.
3.2.1 Merging Captions
In this setup, we create a single caption by merging all the multi-scene descriptions. Specifically, the multi-scene descriptions $Y=\{y_{1},y_{2},\ldots,y_{n}\}$ can be written as a single prompt ‘$\mathcal{P}=y_{1}.\text{Then, }y_{2}.\ldots\text{Then, }y_{n}.$’ For instance, the two-scene description $Y=$‘A red panda climbs on a bamboo forest.’, ‘The red panda sleeps peacefully in the treetop.’} will change to $\mathcal{P}=$ ‘A red panda climbs on a bamboo forest. Then, the red panda sleeps peacefully in the treetop.’ Subsequently, we generate a video from the T2V model $f_{\theta}$ by conditioning it on $\mathcal{P}$. While this approach mentions the temporal sequence of the events in a single prompt, the T2V model does not understand the temporal boundaries between the two events. Specifically, the Eq. 2 suggests that the visual features for all the video frames will aggregate information from the entire multi-scene description, at once, without any knowledge about the alignment between the scene description and its expected appearance in the generated video.
3.2.2 Merging Videos
In this setup, we generate videos for each scene description individually and merge them in the raw input space. Formally, the individual scene description $y_{i}$ conditions the T2V model $f_{\theta}$ to generate the parts of the multi-video $x_{i}$. Finally, we stitch the individual videos together to synthesize the entire video $X={x_{1},x_{2},\ldots,x_{n}}$. In this process, the parts of the multi-scene video closely adhere to the scene descriptions, leading to high text fidelity. However, since the generated videos do not have access to all the multi-scene descriptions (e.g., the video for Scene 2 is not informed about Scene 1), the visual consistency across the entire video is quite poor.
3.3 Multi-Scene Video-Text Data Generation
@@
While our approach generates better multi-scene videos, the text adherence capabilities of the pretrained T2V generative model are limited. This is due to the lack of multi-scene video-text data during its pretraining. Unlike single video-text datasets, the multi-scene video-text datasets are not widely available and are hard to curate for model training. This is attributed to the fact that high-quality caption generation requires a lot of human labor which is time-consuming and expensive. Prior work such as ActivityNet [39] has curated human captions for specific video scenes depicting useful actions in long videos. However, the video scenes are either overlapping or have a large temporal gap between them that will be harmful for natural and smooth variations between the generated multi-scene videos. Hence, the absence of high-quality captions for continuous video scenes in the dataset makes unsuitable for T2V generative training.
@@
To this end, we aim to create a real-world multi-scene video-text dataset to allow further training of the pretrained T2V models. Specifically, we leverage the capability of the multimodal foundation model, Gemini-Pro-Vision [40], to generate high-quality synthetic data for enhanced video-text training [41]. Formally, we start with a video-text dataset $\mathcal{M}=\mathcal{A}\times\mathcal{B}$ consisting of pairs of $(A_{i},B_{i})$ where $A_{i}$ is a raw video and $B_{i}$ is the corresponding video description from the dataset. Subsequently, we utilize PySceneDetect library  [^1] https://github.com/Breakthrough/PySceneDetect to generate continuous video scenes from $A_{i}=\{A_{i,1},A_{i,2},\ldots,A_{i,m}\}$ where $m$ is the number of scene cuts in the video. A similar approach was used in a prior work [12] to detect scene changes in the video data. Then, we sample the middle video frame $F_{i,j}$ as a representative of the semantic content in the video scene $A_{i,j}$. Finally, we input all the video frames $F_{i}=\{F_{i,1},\ldots,F_{i,m}\}$ for a single video $A_{i}$ and the entire video caption $B_{i}$ to a large multimodal model [40]. Specifically, the model is prompted to generate high-quality captions for each of the frames $F_{i,j}$ such they form a coherent narrative guided by the common caption $B_{i}$. We provide the prompt provided to the multimodal model in Appendix §A. In Figure 4 we provide an instance for the multi-scene video-text data generation.
Datasets. To construct a multi-scene video-text dataset, we utilize existing dataset that include natural (real) videos and associated high-quality human-written captions that summarize the entire video. Specifically, we choose MSR-VTT [42] and VaTeX [43]. Most of the videos in MSR-VTT are 10-30 seconds long while VaTeX consists 10 seconds long videos. In addition, each video in MSR-VTT and VaTex consists 20 captions and 10 captions, respectively, out of which one is selected at random for multi-scene data generation. As described above, a single video is cut into multiple video segments using Pyscene library. In our experiments, we retain the first four video segments and discard any additional segments if the library generates more than four. Since the accuracy of the multi-scene captioning and the computational demands during finetuning are influenced by the number of scenes, we opt to limit the scene count to four for our experiments. However, future work could employ similar methodologies to scale the number of scenes, given more computing power and advanced multi-scene captioning models. We provide the data statistics for the final multi-scene data in Appendix §G.
4 Evaluation
In this section, we describe the evaluation scheme for videos generated from multi-scene text descriptions. First, we describe the evaluation metrics that we aim to assess in this work (§4.1). Then, we generate multi-scene descriptions for a diverse set of tasks (§4.2). Finally, we present the details for automatic and human evaluation of the generated videos (§4.3).
4.1 Metrics
The ability to assess the quality of the generated multi-scene videos is a challenging task itself. As humans, we can judge the multi-scene videos across diverse perceptual dimensions [44] that the existing automatic methods often fails to capture [45]. Following [28], we focus on the visual consistency of the generated video, text adherence capabilities of the T2V models, and video quality of the video. Here, we present the metrics with the aspects that they intend to assess in the generated video for multi-scene text description.
Visual Consistency. This metric aims to assess the (entity or background) consistency between the frames of the multi-scene videos. Here, the entity consistency aims to test whether the entities in the multi-scene video are consistent across the video frames. For instance, the appearance of an animal should not change without a change described in the text description. In addition, the background consistency aims to test whether the background of the multi-scene video remains consistent across the video frames. For instance, the room should not change without a change description in the text.
Text Adherence. This metric aims to test whether the generated video adheres to the multi-scene text description. For instance, the events and actions described in the text script should be presented in the video accurately, and in the correct temporal order.
In our experiments, we compute the visual consistency and text adherence with the automatic and human evaluators. Further, we compute the overall score, which is the average of the visual consistency and text adherence scores. In addition, we also assess the visual quality of the generated videos using human evaluation to understand whether the video contains any flimsy frames, shaky images, or undesirable artifacts (Table 1.
4.2 Task Prompts
Here, we curate a set of task prompts for diverse scenarios, aiming to holistically assess the quality of the generated videos.
@@ 4.2 Task Prompts
Single character in multiple visual contexts (S1). In this scenario, we instruct an LLM, GPT-4, to create a coherent script consisting of four scenes. Each scene features a specific animal character performing diverse activities in every scene. This task assesses the capability of the T2V model to generate consistent appearance of the entity and its background while adhering to the different actions (or events) described in the multi-scene text script. For instance, a generated script could be ‘Scene 1: A red panda is climbing a tree. Scene 2: The red panda eats the leaves on the tree. Scene 3: The red panda lies down on the branch of the tree. Scene 4: The red panda sleeps on the branch’. In total, we generate $100$ prompts in this scenario.
@@
Different characters in a specific visual context (S2). In this scenario, we instruct a language model, GPT-4, to create a coherent script consisting of four scenes. Each scene features different animal characters engaging in the same activity in every scene [20]. This task assesses the capability of the T2V model to generate consistent appearance of the background while adhering to the appearance of the different characters in the multi-scene text script. For instance, a generated script could be ‘Scene 1: A cat leaps onto countertop. Scene 2: A dog leaps onto the same countertop. Scene 3: A rabbit leaps onto the same countertop. Scene 4: A raccoon leaps onto the same countertop’. In total, we generate $100$ prompts in this scenario.
@@
Multi-scene captions from real videos (S3). Here, we aim to assess the ability of the model to generate multi-scene videos for open-ended prompts that are derived from real-world videos. This task also assesses the ability of the T2V model to generate consistent appearances of the entity and its background while adhering to multi-scene descriptions. Specifically, we use our multi-scene video-text data generation pipeline (§3.3) to create such prompts for the real videos from the test splits of the video-text datasets. For example, a multi-scene text script could be ‘Scene 1: A beauty vlogger introduces her skincare routine. Scene 2: She applies a serum to her face, smoothing it in’. We present a sample of the various task prompts in the Appendix §B. In total, we generate $100$ prompts in this scenario.
4.3 Evaluator
In this work, we devise an automatic evaluation framework and perform human evaluation to assess the quality of the multi-scene generated videos.
@@ 4.3 Evaluator
Automatic Evaluation. Here, we utilize the capability of a large multimodal model, GPT-4-Vision [46], to reason over multiple image sequences. First, we sample four video frames, uniformly, from each scene in the generated video (e.g., 8 videos frames for two-scene video). Then, we prompt the multimodal model with the temporal sequence of video frames from different scenes and the multi-scene text description. Specifically, we instruct the multimodal model to decide the quality of the generated video across various metrics including entity consistency, background consistency, and text adherence. For each metric, the multimodal model assigns one of three possible response $\{yes=1,partial=0.5,no=0\}$. For instance, $yes$ for the entity consistency metric implies that the video frames sampled from the generated video have consistent appearance of the entity described in the multi-scene script. In this work, we do not utilize any existing video-text alignment models [47, 41] for evaluating text adherence as they are trained on single-scene video-text datasets. We present the automatic evaluation prompt in Appendix §C.
Human Evaluation. We also conduct a human evaluation to assess the multi-scene generated videos along the dimensions of visual consistency, text adherence, and visual quality. Specifically, we ask the annotators from Amazon Mechanical Turk (AMT) to choose one of three options for each metric $\{yes,partial,no\}$, similar to the automatic evaluation. In addition, we choose the annotators that pass a preliminary qualification exam. We present the screenshot of the UI in Appendix §D.
4.4 Evaluation Setup
@@
Since merging captions (§3.2) and TALC3.1) methods input the entire multi-scene text description at once, the quality of the video generated by these methods is influenced by the number of scenes described in the text script. Hence, we calculate the performance of the baselines and TALC by averaging the scores assigned to videos generated for two, three, and four scenes. Additionally, we report on visual consistency by averaging the performance across the entity and background consistency metrics. Here, the entity consistency scores are calculated for the task prompts S1 and S3 (since S2 aims to change the characters across scenes), and the background consistency and text adherence scores are computed for all the task prompts. We also evaluate the impact of TALC-based finetuning on the single scene generation in Appendix §I.
@@ 5 Experiments
5.1 Text-to-Video Generative Models
In this work, we utilize ModelScope [10] and Lumiere [6] T2V models for multi-scene video generation. Here, ModelScope is an open-source T2V model with 1.7 billion parameters including the video encoder, text encoder, and denoising U-net network. Specifically, it is trained to generate 16 video frames on the mix of WebVid [14] video-text dataset and LAION [48] image-text dataset. We perform most of our experiments on ModelScope due to its easy-of-access and adoption in prior works [28]. In addition, we also include Lumiere-T2V, a model that leverages space-time U-Net denoising networks to generate high-quality videos. In this work, we include early experiments with Lumiere to showcase the flexibility of the TALC approach for multi-scene video generation.
Base model with TALC. As described in §3.1, our approach modifies the traditional text-conditioning mechanism to be aware of the alignment between text descriptions and individual video scenes. By design, the TALC framework can be applied to the base T2V model during inference, without any multi-scene finetuning. Thus, we compare the performance of the multi-scene videos generated from ModelScope and Lumiere T2V base models under three settings: merging captions, merging videos, and TALC. In this setting, we generate $16$ frames per scene from ModelScope and $80$ frames per scene from Lumiere. We provide more details on the inference in Appendix §F.
Finetuning with TALC. Since the base model is pretrained with single-scene data, we aim to show the usefulness of TALC framework when we have access to the multi-scene video-text data. To this end, we finetune ModelScope on the multi-scene video-text data (§3.3) with TALC framework. As a pertinent baseline, we also finetune the ModelScope without TALC framework by naively merging the scene-specific captions in the raw text space. In this setting, we finetune the T2V model with $8$ frames per scene and the maximum number of scenes in an instance is set to $4$. We provide further details on the finetuning setup in Appendix §H. The inference settings are identical to the prior method of generating videos from the base model without finetuning.
@@ 5.1 Text-to-Video Generative Models
In this section, we present the results for the baselines and TALC framework averaged over a diverse task prompts and multiple scenes using automatic evaluation (§5.2) and human evaluation (§5.3). Finally, we provide qualitative examples for the multi-scene generated videos to showcase the usefulness of our approach (§5.4).
5.2 Automatic Evaluation
@@
In Figure 5(a), we find that the overall score, average of visual consistency and text adherence, of the multi-scene videos generated using the base ModelScope with TALC ($68.6$ points), outperforms the overall score achieved by the videos generated using merging captions ($61.7$ points) and merging videos ($67.5$ points) with the base ModelScope. Specifically, we observe that the visual consistency of the generated video is high for merging captions ($91$ points) and TALC ($89.9$ points) while it is low for merging videos ($65$ points). This indicates that merging videos independently for the individual scene descriptions does not preserve the background and entity appearances across the different frames. In addition, we observe that the text adherence using TALC outperforms merging captions by $14.8$ points, while the text adherence is the highest with a score of $70$ points using merging videos. This can be attributed to the design of the merging videos baseline where individual video scenes adhere to the scene-specific descriptions well. Hence, merging videos independently approach can be viewed as an upper bound on the text adherence metric.
In Figure 5(b), we observe similar trends for the Lumiere T2V generative model. Specifically, we find that the overall score for TALC outperforms merging captions and merging videos by $4$ points and $2$ points, respectively. In addition, we observe that merging captions and TALC achieve a high visual consistency score while merging videos independently has poor visual consistency. Further, we find that TALC outperforms merging captions by $5$ points on text adherence, while merging videos achieves the highest text adherence $65$ points. This highlights that the model more easily generates multi-scene videos that adhere to individual text scripts, whereas adherence to the text diminishes when the model is given descriptions of multiple scenes all at once.
@@
To perform fine-grained analysis of the performance, we assess the visual consistency and text adherence scores for the baselines and TALC framework across diverse task prompts and number of scenes on ModelScope. We present their results in Appendix §E. In our analysis, we find that finetuning with TALC achieves the highest overall score over the baselines across all the scenarios. In addition, we notice that the highest performance is achieved in the scenario that consist of the different entities in a specific visual context. Further, we observe that the performance of the all the methods reduces when the task prompts get more complex i.e., multi-scene captions from real videos. In addition, we observe that finetuning with TALC achieves the highest overall score over the baselines across all the number of scenes. Specifically, we observe that the performance of the merging captions and TALC framework reduces as the number of scenes being generated increases. Overall, we show that the TALC strikes a good balance between visual consistency and text adherence to generate high-quality multi-scene videos.
5.3 Human Evaluation

### v3

@@ 3.1 TALC Architecture
with, $\tau$ the timestamp in the diffusion modeling setup, which is applied during training as well as inference. We illustrate the framework in Figure 2.
While TALC aims to equip the generative model with the ability to follow all the events in the multi-scene descriptions, the visual consistency is ensured by the temporal modules (attentions and convolution blocks) in the existing denoiser network. By design, our approach can be applied to the pretrained T2V model during inference.
3.2 TALC Training
@@
We note that the text adherence capabilities of the pretrained T2V generative model are limited. This is due to the lack of multi-scene video-text data during pretraining, since these datasets are not widely available and are hard to curate for model training. Prior work, such as ActivityNet [37], curated human captions for specific video scenes depicting useful actions in long videos. However, these scenes are either overlapping or have a large temporal gap between them, which are less suitable for natural and smooth transitions required in multi-scene videos.
@@
To this end, we propose a method for creating synthetic real-world multi-scene video-text examples, with which we further train our T2V model. Formally, we start with a video-text dataset $\mathcal{M}=\mathcal{A}\times\mathcal{B}$ consisting of pairs of $(A_{i},B_{i})$ where $A_{i}$ is a raw video and $B_{i}$ is the corresponding video description. Then, similar to [9], we apply PySceneDetect [^2] https://github.com/Breakthrough/PySceneDetect to segment $A_{i}$ into continuous video scenes $A_{i}=\{A_{i,1},A_{i,2},\ldots,A_{i,m}\}$ where $m$ is the number of detected scenes. Next, we sample the middle video frame $F_{i,j}$ in each scene $A_{i,j}$ as its semantic representative. Finally, we input all the video frames $F_{i}=\{F_{i,1},\ldots,F_{i,m}\}$ for $A_{i}$ and the entire video caption $B_{i}$ to Gemini-Pro-Vision-1.0 [38]. Gemini is prompted to generate captions for each of the frames $F_{i,j}$ such they form a coherent narrative guided by the common caption $B_{i}$. In Figure 3, we provide an instance for the multi-scene video-text data generation. We provide the prompt provided to Gemini and the source datasets in Appendix §D and §E, respectively.
4 Experimental Setting
In this section, we describe the setup of our experiments for generated multi-scene videos, including eval-set creation, metrics, automatic and human evaluation, and the tested models.
4.1 Eval-Set Task Prompts
To create an eval-set for testing models that generate multi-scene videos, we curate a set of task prompts for diverse scenarios (The full task prompts are in Appendix §F.). We generate $100$ prompts for each scenario.
@@ 4.1 Eval-Set Task Prompts
Single character in multiple visual contexts (S1): We instruct GPT-4 [39] to create a coherent script consisting of four scenes. Each scene features a specific animal character performing diverse activities in every scene. This task assesses the consistent appearance of the entity and its background while adhering to the different actions (or events) described in the multi-scene text script. An example script is “Scene 1: A red panda is climbing a tree. Scene 2: The red panda eats the leaves on the tree. Scene 3: … Scene 4: …”.
@@
Different characters in a specific visual context (S2): We instruct GPT-4 to create a coherent script consisting of four scenes. Each scene features the same activity but with different animal characters [18]. This task assesses the capability of the T2V model to generate consistent appearance of the background while adhering to the animal character changes. An example script is “Scene 1: A cat leaps onto a countertop. Scene 2: A dog leaps onto the same countertop. Scene 3: A rabbit … Scene 4: A raccoon …”.
@@
Multi-scene captions from real videos (S3): We use our multi-scene video/text data generation pipeline (§3.2) to create similar prompts for real videos for the test splits of the video-text datasets. This task assesses the ability of a model to generate multi-scene videos with consistent appearances of entities and background for open-ended prompts, which are derived from real videos. An example script is “Scene 1: A vlogger introduces her skincare routine. Scene 2: She applies a serum to her face. Scene 3: …”.
4.2 Evaluation Procedure
Following [27], we assess the capability of the tested T2V models to: (a) follow the text instructions in the script (Text Adherence); (b) keep consistent appearance of entities and background across scenes (Visual Consistency), unless the script specifically requests a change in any of them.
We compute the visual consistency and text adherence with automatic and human evaluators. We also report an overall score, which is the average of the visual consistency and text adherence scores. In addition, we also assess the visual quality of the generated videos using human evaluation to understand whether the video contains any flimsy frames, shaky images, or undesirable artifacts (Appendix M):
@@ 4.2 Evaluation Procedure
Automatic Evaluation: We utilize GPT-4-Vision (GPT4V) [40] to reason over multiple images. First, we sample four video frames, uniformly, from each scene in the generated video (e.g., 8 videos frames for a two-scene video). Then, we prompt GPT4V with the temporal sequence of the sampled frames and the multi-scene text description and instruct GPT4V to decide on the quality of the generated video across several attributes, including entity consistency, background consistency, and text adherence. For each attribute, GPT4V assigns one of three possible response $\{\textit{yes}=1,\textit{partial}=0.5,\textit{no}=0\}$. The automatic evaluation prompt is given in Appendix §G.
[…]
@@
Since merging captions (§4.3.1) and TALC3.1) methods input the entire multi-scene text description at once, the quality of the video generated by these methods is influenced by the number of scenes described in the text script. Hence, we calculate the performance of the tested models by averaging the scores assigned to videos generated for two, three, and four scenes, over 900900900900333100 (examples per scenario) * 3 (scenarios) * 4 (2-scene/3-scene/4-scene) = 900. multi-scene prompts. We report visual consistency by averaging the performance across the entity and background consistency metrics. Here, the entity consistency scores are calculated for the task prompts S1 and S3 (since S2 aims to change the characters across scenes), while the background consistency and text adherence are computed for all the task prompts.
4.3 Tested Models
We experiment with ModelScope [10] and Lumiere [11] T2V models for multi-scene video generation with TALC. We mainly experiment with ModelScope due to its ease of reproduction and adoption in prior work [27]. We also experiment we Lumiere to showcase TALC’s generalization.
By design, TALC can be applied to a base T2V model during inference, without any multi-scene finetuning. We compare the multi-scene performance of the two base models without finetuning under three settings: TALC, merging captions, and merging videos (see baseline details in Section 4.3.1).
In addition, we finetune ModelScope with the TALC framework using the multi-scene video-text data (§3.2). As a baseline, we also finetune ModelScope without the TALC framework by naively merging the scene-specific captions in the raw text space. In both cases, we finetune the T2V model with $8$ frames per scene and the maximum number of scenes in an instance is set to $4$. We provide more finetuning details in Appendix §L.
During inference, we generate $16$ frames per scene from ModelScope and $80$ frames per scene from Lumiere. More inference details are found in Appendix §J.
4.3.1 Baselines
For Merging Captions, we create a single caption by merging all the multi-scene descriptions. Specifically, the multi-scene descriptions $Y=\{y_{1},y_{2},\ldots,y_{n}\}$ can be written as a single prompt ‘$\mathcal{P}=y_{1}.\text{Then, }y_{2}.\ldots\text{Then, }y_{n}.$’. For instance, the two-scene description $Y=$‘A red panda climbs on a bamboo forest.’, ‘The red panda sleeps peacefully in the treetop.’} would change to $\mathcal{P}=$A red panda climbs on a bamboo forest. Then, the red panda sleeps peacefully in the treetop.” Subsequently, we generate a video from the T2V model $f_{\theta}$ by conditioning it on $\mathcal{P}$.
For Merging Videos, we generate a video $x_{i}$ for each scene description $y_{i}$ individually and stitch the output videos together afterwards to synthesize the entire video $X={x_{1},x_{2},\ldots,x_{n}}$.
@@ 5 Experiments
5.1 Automatic Evaluation Results
@@
In Figure 4(a), we find that the overall score, average of visual consistency and text adherence, of the multi-scene videos generated using the base ModelScope with TALC ($69$ points), outperforms the overall score achieved by the videos generated using merging captions ($62$ points) and merging videos ($68$ points) with the base ModelScope. Specifically, we observe that the visual consistency of the generated video is high for merging captions ($91$ points) and TALC ($89.9$ points) while it is low for merging videos ($65$ points). This indicates that merging videos independently for the individual scene descriptions does not preserve the background and entity appearances across the different frames.
Since merging captions mentions multiple events in a single prompt, the T2V model struggles to understand the temporal boundaries between these events. As indicated by Eq. 2, the visual features for all video frames aggregate information from the entire multi-scene description simultaneously, without aligning individual scene descriptions with their corresponding appearances in the generated video. In Figure 4(a), this results in poor text adherence with merging captions ($32$ points).
In addition, we observe that the text adherence using TALC outperforms merging captions by $14.8$ points, while the text adherence is the highest with a score of $70$ points using merging videos. This can be attributed to the design of the merging videos baseline where individual video scenes adhere to the scene-specific descriptions well. Hence, merging videos independently approach can be viewed as an upper bound on the text adherence metric.
In Figure 4(b), we observe similar trends for the Lumiere T2V generative model. Specifically, we find that the overall score for TALC outperforms merging captions and merging videos by $4$ points and $2$ points, respectively. In addition, we observe that merging captions and TALC achieve a high visual consistency score while merging videos independently has poor visual consistency. Further, we find that TALC outperforms merging captions by $5$ points on text adherence, while merging videos achieves the highest text adherence $65$ points. This highlights that the TALC framework enhances the adherence to the scene descriptions.
@@
To perform fine-grained analysis of the performance, we assess the visual consistency and text adherence scores for the baselines and TALC framework across diverse task prompts and number of scenes on ModelScope. We present their results in Appendix §I. In our analysis, we find that finetuning with TALC achieves the highest overall score over the baselines across all the scenarios. In addition, we notice that the highest performance is achieved in the scenario that consist of the different entities in a specific visual context. Further, we observe that the performance of the all the methods reduces when the task prompts get more complex i.e., multi-scene captions from real videos. In addition, we observe that finetuning with TALC achieves the highest overall score over the baselines across all the number of scenes. Specifically, we observe that the performance of the merging captions and TALC framework reduces as the number of scenes being generated increases. Overall, we show that the TALC strikes a good balance between visual consistency and text adherence to generate high-quality multi-scene videos.
5.2 Human Evaluation Results